The demand for ice cream cones is P=1600 and Qd is 2
The supply of ice cream cones is P =400 and Qs is 1. The price of a cone is expressed in cents, and the quantities are expressed in cones per day. To find the equilibrium price (P*) and the equilibrium quantity (Q*), substitute Q* for QD and QS and P* for P. (8 marks). (FOR QD, P=a-bQd and for QS, P=c+dQS).
1
Expert's answer
2019-04-05T10:36:08-0400

Demand function is P=a - bQd  

Supply function is P=c+dQs, so:

1600 = a - 2b, 400 = c + d,

In equilibrium Qd = Qs

a – bQ* = c + dQ*,

a – c = bQ* + dQ*

Q* = (a - c)/(b + d).